OpenStudy (anonymous):

I'll give medals!!!!! The sum of three consecutive numbers is 57. What is the largest of these numbers? 18 19 20 21

OpenStudy (anonymous):

ive got it!! is it 20

OpenStudy (shinalcantara):

Let 'x' be the 1st number 'x+1' be the 2nd number 'x+2' be the 3rd number ------------------------- Since the sum of the 3 is 57 then x + x+1 + x+2 = 57 x + x + x = 57 - 1 - 2 3x = 54 x = 18 x+1 =19 x+2=20

OpenStudy (anonymous):

once again thanks @shinalcantara